Solve Problems Using Linear Systems
14. Mengxi has $10 000 to invest. She invests part in a term deposit paying 5%/year, and the remainder in Canada Savings Bonds paying 3.5%/year. At the end of the year, she has earned simple interest of $413. How much did she invest at each rate?
Thank you very much>>

Answer by badbadjoe(1)

.05x+.035(10,000-x)=413
.05x+350-.035x=413
.015x+350=413
.015x=63
x=4200
In the 5% account she invested $4200 and
in the 3.5% account she invested $5800